OpenStudy (mtalhahassan2):

Milk is leaking from a carton at a rate of 3mL/min. There is 1.2 L OF MILK IN THE CARTON AT 11:00A.a.m. a) Use function notation to write an equation for this situation. b) How much will be left in the carton at 1:00p.m.? c) At what time will 450 mL of milk be left in the carton?

OpenStudy (ash2326):

At 11:01, 3mL of Milk would have been leaked from the carton, so the remaining milk at 11:01 am would be M(1)=1200 mL-3mL M(1) denotes the Milk left after 1 minute has passed. Do you understand?

OpenStudy (mtalhahassan2):

no

OpenStudy (ash2326):

Where do you have doubt?

OpenStudy (anonymous):

@ash2326 You should have shown how you substituted the 1 into the function. M(x)=1200-3x M(1)=1200-3(1)=1200-3

OpenStudy (ash2326):

@jedi_trooper Yes, I was going to come to that part. By showing that we can create function in terms of x minutes after 11 am.

OpenStudy (anonymous):

I think you should have started with that. For a person who has not seen functions before, it can be confusing to see M(1). They'll wonder where did the 1 come from or why do you use 1 instead of 2, etc.

OpenStudy (mtalhahassan2):

so what be the answer of part c

OpenStudy (anonymous):

450=1200-3x. Solve for x. X is in minutes so add that amount of minutes to your starting time to get your ending time.